Singly charged sodium atoms are accelerated through a potential difference of 342 V. (a) What is the momentum acquired by such an ion? (b) What is its de Broglie wavelength? The mass of a sodium ion is 3.819 × 10-26 kg.
Answer #1

(A) kinetic energy = qV

0.5mv^2 = qV

0.5*3.819*10^-26*v^2 = 1.6*10^-19*342

v = 5.35*10^4 m/s

momentum = mv = 3.819*10^-26*5.35*10^4

             = 2*10^-21 kg m/s

(b) E = hc / \lambda

1.6*10^-19*342 = 6.626*10^-34*3*10^8 / \lambda

\lambda = 3.63 *10^-9 m

   = 3.63 nm